the rocking pebbles are playing a concert at big state university the auditoruim seats 2200 people.the concert manager decides to sell some tickets at $10 each and the rest at $15 each. how many of each kind should the manager sell if he wants ticket sales to total $26,600? how do i do this
Answer by mananth(16949)

$10 tickets x
$15 tickets y
1 x + 1 y = 2200 .............1
10 0 + 15 y = 26600 .............2
1 x + 1 y = 2200 .............1
10 x + 15 y = 26600 .............2
Eliminate y
multiply (1)by -15
Multiply (2) by 1
-15x-15y=-33000
10x+15y=26600
Add the two equations
-5x=-6400
/-5
x=1280 tickets
plug value of x in (1)
x+y=2200
1280+y=2200
y = 2200 -1280
y = 920
y = 920 tickets